# Embeddable partial groups

### v1

@@ Abstract.
This work was supported by a grant from the Simons Foundation (#850849, PH) and by a grant from the Simons Foundation International (SFI-MPS-TSM-00014188, JL). PH was partially supported Louisiana Board of Regents through the Board of Regents Support fund LEQSF(2024-27)-RD-A-31.
@@ 1. Embeddability
Theorem˜2 and Corollary˜3 are many-object versions of Corollary 1 and Corollary 2 of [3, §4]. Indeed, a 2-skeletal partial group [14] is essentially the same thing as what Baer termed an ‘add which is self-reflexive in the strict sense’: in the forward direction one sends $X$ to $X_{1}$ equipped with its partial binary multiplication, while in the reverse direction one builds a partial group whose set of 2-simplices is the set of multipliable pairs. Baer’s theorems were later rediscovered by Tamari [21, §III], who called these partial magmas monoïdes symétriques. The one-object version of Theorem˜2 was later known as Tamari’s One Mountain Theorem, e.g. [4, Theorem 2.a].
@@
Fundamental groups of linking systems of saturated fusion systems are of interest in $p$-local group theory [2, §III.7 (8)]. Linking systems are special kinds of transporter systems [19, §3], which are equivalent to special partial groups called localities [5], [9, §2]. The locality $L$ associated with a transporter system $\mathcal{T}$ is a quotient $N\mathcal{T}\twoheadrightarrow L$ of the nerve, and this quotient map is a weak equivalence [10, Appendix]. Theorem˜2 may thus be interpreted in light of the computations of fundamental groups of linking systems that have already been carried out. As one example, $\pi_{1}(L^{c}_{\mathrm{Sol}}(q))=1$ for the centric linking locality of a Benson–Solomon fusion system [6]. By Theorem˜2 any two elements of $L^{c}_{\mathrm{Sol}}(q)$ have some word associating to each.
@@ 2. Orthogonality set
Suppose $(i-1,i,i+1)\notin T\cap T^{\prime}$ for any $i$. To prove that the 2-dimensional symmetric set $X$ is spiny, it suffices to show that $\mathscr{E}_{2}\colon X_{2}\to X_{1}\times_{X_{0}}X_{1}$ is injective [14, Theorem 9]. As the subobjects $H(T)$, $H(T^{\prime})$ of $X$ are spiny, it suffices to consider $x\in H(T)_{2}$ and $x^{\prime}\in H(T^{\prime})_{2}$ such that $\mathscr{E}_{2}(x)=\mathscr{E}_{2}(x^{\prime})$ in $X$. Since the intersection of $H(T)$ and $H(T^{\prime})$ in $X$ is $\textup{Sp}^{n}$, both edges
@@ 4. Reduction and embeddability
For convenience, we rephrase the definition of this monoid in terms of rewrite systems and confluence [16, §2.1–2.2]. Set $S$ be the set of words in $C_{1}$, and consider the following rewrite rule $\rightarrow$ on $S$:
@@
Suppose $X$ is embeddable. By factoring the composite $X\to N\tau X\to\mathcal{R}N\tau X$ through $\mathcal{R}X$ and applying Lemma˜23 to the category $\tau X$, we have the following diagram
@@
The reduction of an injective map is injective, so $\mathcal{R}X$ embeds in the group $M(\tau X)$. The converse is Lemma˜19. ∎

### v2

@@ Abstract.
This work was supported by a grant from the Simons Foundation (#850849, PH) and by a grant from the Simons Foundation International (SFI-MPS-TSM-00014188, JL). PH was partially supported Louisiana Board of Regents through the Board of Regents Support fund LEQSF(2024-27)-RD-A-31. ES is grateful for travel support from TU Nachwuchsring at the RPTU Kaiserslautern-Landau.
@@ 1. Embeddability
Theorem˜2 and Corollary˜3 are many-object versions of Corollary 1 and Corollary 2 of [3, §4]. Indeed, a 2-skeletal partial group [15] is essentially the same thing as what Baer termed an ‘add which is self-reflexive in the strict sense’: in the forward direction one sends $X$ to $X_{1}$ equipped with its partial binary multiplication, while in the reverse direction one builds a partial group whose set of 2-simplices is the set of multipliable pairs (for details see [14]). Baer’s theorems were later rediscovered by Tamari [22, §III], who called these partial magmas monoïdes symétriques. The one-object version of Theorem˜2 was later known as Tamari’s One Mountain Theorem, e.g. [4, Theorem 2.a].
@@
Fundamental groups of linking systems of saturated fusion systems are of interest in $p$-local group theory [2, §III.7 (8)]. Linking systems are special kinds of transporter systems [20, §3], which are equivalent to special partial groups called localities [5], [9, §2]. The locality $L$ associated with a transporter system $\mathcal{T}$ is a quotient $N\mathcal{T}\twoheadrightarrow L$ of the nerve, and this quotient map is a weak equivalence [10, Appendix]. Theorem˜2 may thus be interpreted in light of the computations of fundamental groups of linking systems that have already been carried out. As one example, $\pi_{1}(L^{c}_{\mathrm{Sol}}(q))=1$ for the centric linking locality of a Benson–Solomon fusion system [6]. By Theorem˜2 any two elements of $L^{c}_{\mathrm{Sol}}(q)$ have some word associating to each.
@@ 2. Orthogonality set
Suppose $(i-1,i,i+1)\notin T\cap T^{\prime}$ for any $i$. To prove that the 2-dimensional symmetric set $X$ is spiny, it suffices to show that $\mathscr{E}_{2}\colon X_{2}\to X_{1}\times_{X_{0}}X_{1}$ is injective [15, Theorem 9]. As the subobjects $H(T)$, $H(T^{\prime})$ of $X$ are spiny, it suffices to consider $x\in H(T)_{2}$ and $x^{\prime}\in H(T^{\prime})_{2}$ such that $\mathscr{E}_{2}(x)=\mathscr{E}_{2}(x^{\prime})$ in $X$. Since the intersection of $H(T)$ and $H(T^{\prime})$ in $X$ is $\textup{Sp}^{n}$, both edges
@@ 3. Degree of $\operatorname{NA}^{T,T^{\prime}}$
Below we give an alternative characterization of degree in terms of the ternary associativity condition mentioned in Example˜8: a partial groupoid $X$ satisfies $\diamondsuit$ if its underlying partial binary composition has the property that whenever $ab$ and $bc$ are defined, then $(ab)c$ is defined if and only if $a(bc)$ is defined, with equality when both hold.
Lemma 19.
If a partial groupoid $X$ does not satisfy $\diamondsuit$, then $\deg(X)>2$.
Proof.
If $X$ does not satisfy $\diamondsuit$, then there is a word $w=(c,b,a)\in\mathbf{W}(X)_{3}$ such that $ab$, bcbc, and $a(bc)$ exist, but either $(ab)c$ does not exist or it does and $(ab)c\neq a(bc)$. The faces $d_{0}w=[b|a]$, $d_{1}w=[bc|a]$, and $d_{3}w=[c|b]$ are in $X_{2}$, but $w\notin X_{3}$. Thus $X$ is not lower 3-Segal by [12, Remark 3.17]. ∎
If $G$ is a group with at least four elements, then the 2-skeleton of $BG$ satisfies $\diamondsuit$ and has degree 3, so the converse of the lemma does not always hold.
Proposition 20.
Let $(T,T^{\prime})$ be a compatible pair of triangulations of a regular $(n+1)$-gon and $X=\operatorname{NA}^{T,T^{\prime}}$. Then $X$ satisfies $\diamondsuit$ if and only if $\deg(X)=2$.
Proof.
The backwards direction follows immediately from Lemma˜19. We use Proposition˜18 for the forward direction: if $X$ has degree 3, then $(T,T^{\prime})$ has a cone, and without loss of generality we assume there are triangles (i1,i,k),(i,i+1,k)(i-1,i,k),(i,i+1,k) in $T$ and (i1,i,i+1)(i-1,i,i+1) in $T^{\prime}$. There are edges $c\colon i-1\to i$ and $b\colon i\to i+1$ in $\textup{Sp}^{n}$, $a\colon i+1\to k$, $ab\colon i\to k$, and $(ab)c\colon i-1\to k$ in $H(T)$, and $bc\colon i-1\to i+1$ in $H(T^{\prime})$. If $a(bc)$ does not exist, then $\diamondsuit$ does not hold. Suppose $a(bc)$ exists. As $bc$ is not in $\textup{Sp}^{n}$, hence not in $H(T)$, it must be the case that $a\in\textup{Sp}^{n}$. Since $k\neq i,i+1$, we have $k=i+2$ and (i1,i+1,i+2)(i-1,i+1,i+2) is a triangle of $T^{\prime}$. But in this case $a(bc)\colon i-1\to i+2$ in $H(T^{\prime})$ is necessarily different from $(ab)c\colon i-1\to i+2$ in $H(T)$, so again $\diamondsuit$ fails. ∎
@@ 4. Reduction and embeddability
For convenience, we rephrase the definition of this monoid in terms of rewrite systems and confluence [17, §2.1–2.2]. Set $S$ be the set of words in $C_{1}$, and consider the following rewrite rule $\rightarrow$ on $S$:
@@
Suppose $X$ is embeddable. By factoring the composite $X\to N\tau X\to\mathcal{R}N\tau X$ through $\mathcal{R}X$ and applying Lemma˜25 to the category $\tau X$, we have the following diagram
@@
The reduction of an injective map is injective, so $\mathcal{R}X$ embeds in the group $M(\tau X)$. The converse is Lemma˜21. ∎